OpenStudy (anonymous):

the perimeter of a certain rectangle is no more than 94 feet if the length is eleven feet more than the width what is the width

OpenStudy (anonymous):

2w + 2(w+11) = 94 solve for w

OpenStudy (anonymous):

4w + 22 = 94 w = 72/4 w = 18